OpenStudy (anonymous):

CAN ANYONE HELP ME SOLVE THESE PATTERNS? 5, 1, 7, 0, 9, 1, 11, . . . 4.5, 0.5, 5.0, 0.6, 5.6, 0.7, 6.3, . . . They are two different questions but i'm kinda stuck.

OpenStudy (akashdeepdeb):

Can you rewrite the patterns?

OpenStudy (akashdeepdeb):

4.5, 0.5, 5.0, 0.6, 5.6, 0.7, 6.3, . . . Here's the second pattern: 4.5 + 0.5 = 5.0 5.0 + 0.6 = 5.6 5.6 + 0.7 = 6.3 6.3 + 0.8 = ____ ____ + 0.9 = .....

OpenStudy (anonymous):

okay thanks